OpenStudy (anonymous):

can any1 solve this? 6 − |3x + 9| = 3

OpenStudy (anonymous):

\[6-|3x+9|=3\]subtract 6 \[-|3x+9|=-3\]change the sign \[|3x+9|=3\] then solve \[3x+9=3\]and\[3x+9=-3\] separately to get two different answers

OpenStudy (anonymous):

i got -2 and -4 check plz

OpenStudy (jiteshmeghwal9):

since,|a|=-a & a . so, -(3x+9)=-3x-9 & 3x+9 put these both values in the equation (i)6-(-3x-9)=3 6+3x+9=3 3x+15=3 3x=12 x=12/3=4 (ii) 6-(3x+9)=3 6-3x-9=3 -3x-3=3 -3x=9 x=9/-3=-3

OpenStudy (jiteshmeghwal9):

hence , answers are 4 & -3